Chris has invested $2,3000 in two small buisnesses. Last year, the greeting card buisness made a 14% profit on her investment in the buisness, while the bicycle repair buiness showed a 6% loss. If Chris' net profit fromm the two buisnesses was$102, her investment in the greeting card buisness was?
Answer by checkley77(12844)

LETS ASSUME THE INVESTMENT IS 23,000
.14X-.06(23,000-X)=102
.14X-1,380+.06X=102
.2X=102+1,380
.2X=1,482
X=1,482/.2
X=$7,410 ANSWER FOR THE AMOUNT INVESTED IN THE CARD BUSINESS.
23,000-7,410=15,590 AMOUNT INVESTED IN THE BIKE SHOP.
PROOF
.14*7,410-.06*15,590=210
1037.4-935.4=102
102=102